1. move the constant term to the right side of the equation
2.muliply each term in the equation by 4 times the coefficient of the x^2
3. square the coefficient of the original x term and add to both sides of the equation
4.take the square root of both sides
5.set the left side of the equation equal to the positive square root of the number on the right side and solve for x
6. set the left side of the equation equal to the negative square root of the number on the right side of the equation and solve for x.


x^2 -2x -13 = 0

x^2 + 12x -64 = 0

I need these two problems worked out so I can learn from them and learn how to do them for my assignment.

Answer by Theo(13342)

let's follow the method using the equation of x^2 - 2x - 13 = 0 and see how it works.
the method is:
1.   move the constant term to the right side of the equation.

     you get x^2 - 2x = 13

2.   multiply each term in the equation by 4 times the coefficient of the x^2

     you get 4 * (x^2 - 2x = 13) becomes 4x^2 - 8x = 52

3.   square the coefficient of the original x term and add to both sides of the equation

     the original coefficient of the x term i 2 so we square it and we get 
     4 and we add it to both sides of the equation to get:
     4x^2 - 8x + 4 = 52 + 4 which becomes:
     4x^2 - 8x + 4 = 56

4.   take the square root of both sides


     the square root of 56 is sqrt(56).


     the square root of 4x^2 - 8x + 4 should be (2x-2) because (2x-2)^2 will
     be equal to 4x^2 - 8x + 4.  
     the equation becomes:
     (2x+2) = +/- sqrt(56)

5.   set the left side of the equation equal to the positive square root of 
     the number on the right side and solve for x

     2x + 2 = sqrt(56) which leads to 2x = -2 + sqrt(56) which leads to
     x = (-2 + sqrt(56)) / 2

6.   set the left side of the equation equal to the negative square root of
     the number on the right side of the equation and solve for x.

     2x + 2 = - sqrt(56) which leads to 2x = -2 - sqrt(56) which leads to
     x = (-2 + sqrt(56)) / 2


what they have showed you is the indian method for solving a quadratic equation.
you should be able to get the same answer using the traditional method.
the original equation is:
x^2 -2x -13 = 0
this is in standard form of ax^2 + bx + c = 0
this leads to:
a = 1
b = -2
c = -13
use the quadratic formula to solve this.
the solution you get from using the quadratic formula is:
x = (-2 +/- sqrt(56)) divided by 2.
this is the same answer we got above using the indian method.
the method is set up in such a way that when they tell you to get the square root of both sides of the equation, you can get the square root of the modified quadratic equation on the left side of the equation.
that modified equation was 4x^2 - 8x + 4 which allowed you to get the square root of it.

you should be able to do the second equation, but if you have trouble with it let me know and i'll help you solve it.

not now though - i have to go.
bye